how do I solve using the substituion method with the following equations. 3b-2a=4 and b/2-2a/3=1?

Answer by mananth(16949)

3b-2a=4
3b= 4+2a
3b=2(2+a)
b=2(2+a)/3
....
and b/2-2a/3=1
Plug value of b in the baove equation
2(2+a)/3*2 - 2a/3 =1
(2+a)/3 -2a/3 =1
...
(2+a-2a)/3 =1
(2-a) /3=1
2-a=3
a=2-3
a=-1
plug value of a in the equation
b=2(2+a)/3
b=2(2-1)/3
b=2/3
...
m.ananth@hotmail.ca